OpenStudy (swamijatin2):

divide 243 into3 parts such that half of the first part ,one-third of the second part and one fourth of the third part are all equal. please tell me..

Directrix (directrix):

@swamijatin2 I transformed the words to variables along with the given relationship and got the following equations: x+y+z =243 .5x = (1/3)y .25z =.5x I used the "Wolf" to solve these 3 equations in three unknowns. The "Wolf" cranked out: x = 54, y = 81, and z = 108 You'll need to determine if all the required relationships of the three numbers to each other hold up. http://www.wolframalpha.com/input/?i=x%2By%2Bz+%3D243+and+.5x+%3D+%281%2F3%29y+and+.25z+%3D.5x

OpenStudy (swamijatin2):

oh i got it thanxs

Directrix (directrix):

Okay